Express sin^5 x as a linear combination of sin kx and coskx, k belongs to Z
1
Expert's answer
2020-03-19T16:59:30-0400

Express sin5 (x) as a linear combination of sin (kx) and cos (kx)

sin5 (x) = sin3 (x) * sin2 (x)


We use the degree reduction formula:

sin2 (x) = (1 - cos (2x))/2

sin3 (x) = (3*sin (x) − sin (3x))/4


We get:

sin5 (x) = (3*sin (x) − sin (3x)) * (1 − cos (2x)) / 8 =

= (3*sin (x) − sin (3x) - 3*sin (x) * cos (2x) + sin (3x) * cos (2x))/8


We use the formula for the product of sines and cosines:

sin (a) * cos (b) = (sin (a+b) + sin (a-b))/2


sin(-a) = - sin(a)


We get:

sin5 (x) = (6*sin (x) − 2*sin (3x) - 3*sin (3x) + 3*sin (x) + sin (5x) + sin (x))/16 =

= (sin (5x) - 5*sin (3x) + 10*sin (x))/16


Answer: sin5 (x) = (sin (5x) - 5*sin (3x) + 10*sin (x))/16